OpenStudy (anonymous):

Simplify e^2.2 .e^3.1

OpenStudy (anonymous):

i assume they are multiplied?

OpenStudy (anonymous):

you can use the rule \[\large x^p*x^q=x^{ p+q} \]

OpenStudy (anonymous):

I think they are multiplied. That was how the question was laid out to me. I wasn't sure if they the letter e had any relation to the ln button on my calculator?

OpenStudy (anonymous):

if there is no sign between them (or just a flying dot) that means multiply.

OpenStudy (anonymous):

the letter e does, but even so, the rule still stands even for e.

OpenStudy (anonymous):

ln is the log with base e.

OpenStudy (anonymous):

\[\Large \ln (e^x) =x\\\Large e^{ \ln(x)} =x\]

OpenStudy (anonymous):

oh, okay.